OpenStudy (anonymous):

Can someone help me find the square root of 17x^2 - 90x + 125?

OpenStudy (anonymous):

did you try the quadratic formula?

OpenStudy (anonymous):

no I will do that now :)

OpenStudy (anonymous):

17=a -90=b 125=c for the quadratic formula

OpenStudy (anonymous):

thanks I already plugged those in, I'm solving now

OpenStudy (anonymous):

I got 90 sqrt -11.764 hows that look? @dariogpx

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

i got 100/34 and -70/34

OpenStudy (anonymous):

im going to check again

OpenStudy (anonymous):

ok i was wrong i am left with a negative 400 in the radical............are you with me @bugzkilla ?

OpenStudy (anonymous):

yeah I'm here I'm trying to attach my work for you

OpenStudy (mrnood):

you asked in your question about 'square root' of the equation - did you mean the 'roots'?

OpenStudy (anonymous):

the part that was cut off was / 34

OpenStudy (anonymous):

the two answers at the end of using the quadratic formula @MrNood

OpenStudy (mrnood):

Be aware that the 'roots' of a quadratic are not related to 'square roots'

OpenStudy (anonymous):

you have to get the negative out of the radical with the imaginary number i so it will be the imaginary number "i" the square root of 400

OpenStudy (anonymous):

i am aware of that i just know what he means @MrNood

OpenStudy (mrnood):

@dariogpx - I am really addressing my point to @bugzkilla HE is the one who wrote 'square root'

OpenStudy (anonymous):

90 \[\sqrt{-90^2 -4(17)(125)}\]

OpenStudy (anonymous):

all divided by 2(17)

OpenStudy (mrnood):

That is correct - but as dariogpx says - you cannot have a real root of a negative number - so the roots are complex in this case.

OpenStudy (anonymous):

\[90\sqrt{-11.764}\]

OpenStudy (anonymous):

ok thanks for the help @MrNood @dariogpx